OpenStudy (anonymous):

AB = 4, BC = 17, CA =

OpenStudy (lgbasallote):

AC = AB + BC try :D

OpenStudy (mathteacher1729):

Something tells me there is more to this story...

OpenStudy (lgbasallote):

CA is a negative of that btw...because it's backwards

OpenStudy (callisto):

Is there a diagram? It is a triangle?

OpenStudy (mathteacher1729):

My spider senses tell me that the goal is to find reasonable values for the third length of a triangle with verticies A,B, and C.

OpenStudy (callisto):

@mathteacher1729 Agree...

OpenStudy (lgbasallote):

or maybe we're just being paranoid and it's really that simple :P

OpenStudy (anonymous):

yes, mathteacher I think thats true

OpenStudy (anonymous):

What is it asking this person to do exactly?

OpenStudy (lgbasallote):

would be nice if he told us huh :))

OpenStudy (callisto):

@keno Can you tell us more about the question?

OpenStudy (anonymous):

A triangle has the following measurements. What is a possible length for the third side AB = 4, BC = 17, CA =

OpenStudy (lgbasallote):

haha you guys were right :P

OpenStudy (lgbasallote):

 \[CA = \sqrt{(17^{2} - 4^{2})}\]

OpenStudy (callisto):

Assume it is a right angle triangle. Use Pyth Thm to solve Case 1, AC is the longest side AB^2 + BC^2 = CA^2 4^2 + 17^2 = CA^2 CA = sqrt (305) Case 2, BC is the longest side AB^2 + CA^2 = BC^2 4^2 + CA^2 = 17^2 CA = sqrt (273)